Please help me solve this equation: 3d- 11(on 4) = 3(on2)
thank you.

Answer by checkley71(8403)

3D-11/4=3/2
3D=3/2+11/4
3D=6/4+11/4
3D=17/4
D=(17/4)/3
D=(17/4)*1/3
D=17/12
PROOF
3*17/12-11/4=3/2
51/12-11/4=3/2
51/12-33/12=18/12
(51-33)/12=18/12
18/12=18/12